Evaluate each expression. Show all your work.
|3+8|
|3-8|
|-3+8|
|3| + |8|

|-3| + |8|

|3| - |8|

Answer by Edwin McCravy(20060)

Evaluate each expression. Show all your work.
Rule for removing absolute values:

1.  Complete all operations, if any, between the absolute values
2.  (a)  If the number between the absolute value bars is positive, merely
    take away the absolute value bars.
    (b)  If the number between the absolute value bars is negative, change
    the sign to positive and take away the absolute value bars
3.  Do the remaining operations, if any. 
 


|3+8| = |11| = 11


|3-8| = |-5| = 5


|-3+8| = |5|


|3| + |8| = 3 + 8 = 11


|-3| + |8| = 3 + 8 = 11


|3| - |8| = 3 - 8 = -5

Edwin